# Towards Stable Machine Learning Model Retraining via Slowly Varying Sequences

### v4

@@ Abstract
We consider the task of retraining machine learning (ML) models when new batches of data become available. Existing methods focus largely on greedy approaches to find the best-performing model for each batch, without considering the stability of the model’s structure across retraining iterations. In this study, we propose a methodology for finding sequences of ML models that are stable across retraining iterations. We develop a mixed-integer optimization formulation that is guaranteed to recover Pareto optimal models (in terms of the predictive power-stability trade-off) and an efficient polynomial-time algorithm that performs well in practice. We focus on retaining consistent analytical insights — which is important to model interpretability, ease of implementation, and fostering trust with users — by using custom-defined distance metrics that can be directly incorporated into the optimization problem. Our method shows stronger stability than greedily trained models with a small, controllable sacrifice in predictive power, as evidenced through a real-world case study in a major hospital system in Connecticut.
@@ 1 Introduction
The adoption of ML models in high-stakes decision spaces such as healthcare [1] increases the need for interpretability and model understanding [2, 3, 4]. Expert users develop trust in AI-driven decision support systems based on both the systems’ alignment with the users’ personal expertise [5, 6] and the systems’ historical performance [7]. Importantly, trust is built over time and requires deep understanding of the underlying model and the resulting analytical insights [8]. In real-world production pipelines, ML models need to be frequently updated as more data becomes available [9]. Significant changes in the model’s structure and the resulting analytical insights [10] may lead to skepticism and hesitation for adoption. Especially in healthcare, such changes may challenge both physicians’ trust in AI systems and patients’ confidence in medical decision-making [11, 12], and may even lead to legal difficulties as a result of incidents of error [13].
Multiple avenues have been proposed in the literature to address challenges associated with model retraining, including continual learning [14, 15], online learning [16], and others [17, 18, 19, 20]. Existing methods focus largely on maintaining the predictive power of the retrained models in the face of, e.g., data distributional shifts. Such methods do not, however, ensure that the decision path of a tree-based prediction or the important features selected by a model remain stable across different retraining iterations, and therefore do not guarantee retention of analytical insights. In this work, we explicitly incorporate this requirement into the (re)training process.
Retraining (batch learning) differs from online learning: the former periodically retrains the model from scratch with each new data batch; the latter continuously/incrementally updates the model as new data arrives. Thus, a batch learning-based retraining strategy allows for a globally optimal solution, as retraining uses the entire database, whereas online learning, often to ensure faster updates, incrementally incorporates only a portion of the data, hence providing greedier solutions [21].
The problem of developing stable ML models is not new in the statistics, computer science, and operations research literature [22, 10, 23]. Various methodologies have recently been proposed to either train structurally stable, interpretable models [24, 25, 26, 27], or black-box models that produce stable feature interactions [28, 29]. In this paper, we develop a unifying, model-agnostic framework for training sequences of models that remain stable upon retraining with new data.
We summarize the main contributions of this paper: (i) We develop a general framework aimed at stabilizing ML model structures upon retraining with new data. We propose an exact mixed-integer optimization (MIO) formulation as well as a heuristic, restricted formulation for the problem. Using the latter, we develop a practical, model-agnostic algorithm for retraining ML models. We also provide theoretical guarantees on the Pareto optimality and generalization of the solutions obtained by our methodology. (ii) We demonstrate across a variety of models (including logistic regression, classification trees, and boosted trees) that our algorithm gives a significant boost in model stability with a small and controllable sacrifice of predictive power and computational efficiency compared to existing greedy approaches. (iii) We show our method’s applicability in an existing production pipeline used in a large US hospital for mortality prediction.
2 A Methodology for Retraining Machine Learning Models
@@ 2.1 Problem Setting
Consider an observational dataset accumulated over time composed of triplets $(\bm{x_{i}},y_{i},t_{i})_{i=1}^{N}$, where $\bm{x_{i}}\in\mathcal{X}\subseteq\mathbb{R}^{p}$ represents the feature vector, $y_{i}\in\mathcal{Y}\subseteq\mathbb{R}$ is the target variable, and $t_{i}\in\mathbb{R}$ is the time at which data point $i$ was obtained. The dataset is divided into a series of batches at times $T_{1}<T_{2}<\ldots<T_{B}$. The dataset available at time $T_{b},\ b\in[B]:=\{1,\ldots,B\},$ is defined as $D_{b}=\{(x_{i},y_{i}):t_{i}\leq T_{b},i\in[N]\}.$ Thus, we have that $D_{1}\subseteq D_{2}\subseteq\ldots\subseteq D_{B}$ and $\bm{D}=(D_{1},\dots,D_{B})$. We assume that $B\geq 2$ and that, at model training time (denoted by $t$), all $B$ batches are available, i.e., $t\geq T_{B}$. In Section 2.5, we discuss how to adaptively retrain the model using batches obtained after time $t$. The goal is to find models $\bm{f}:=(f_{1},f_{2},\ldots,f_{B})\in\mathcal{F}\times...\times\mathcal{F}:=\mathcal{F}^{B}$ with $f_{b}:\mathcal{X}\mapsto\mathcal{Y}$ for $b\in[B]$ such that the following two objectives are satisfied:
Predictive power: The first objective is to ensure that each model $f_{b}$ in the sequence has small expected loss in terms of prediction quality. We denote by $\mathbb{P}(D_{b})$ the distribution of the training data in $D_{b}$ and $\ell:\mathcal{Y}\times\mathcal{Y}\mapsto\mathbb{R}_{\geq 0}$ the bounded loss function measuring, e.g., regression or classification accuracy; concretely, $\ell(y,f_{b}(\bm{x}))$ quantifies the quality of model $f_{b}$’s prediction on input $\bm{x}$ for $y$. Then, we seek for a sequence whereby each model (individually) achieves low expected loss $\mathcal{L}_{P}(f_{b},D_{b})=\mathbb{E}_{(\bm{x},y)\sim\mathbb{P}(D_{b})}[\ell(y,f_{b}(\bm{x}))].$ The overall loss of the sequence is then denoted by $\mathcal{L}_{P}(\bm{f},\bm{D})=\sum_{b=1}^{B}\mathcal{L}_{P}(f_{b},D_{b}).$ In practice, we approximate the expected loss using the empirical loss ^$\hat{\mathcal{L}}_{P}(f_{b},D_{b})=\frac{1}{N_{b}}\sum_{i:t_{i}\leq T_{b}}[\ell(y_{i},f_{b}(\bm{x}_{i}))]$. The next assumption bounds the approximation quality as function of the sample size, the function class complexity, and the desired confidence:
Assumption 1.
@@
Stability: The second objective is to ensure that consecutive models do not vary significantly, maintaining stability over time. Formally, we define a stability metric $d:\mathcal{F}\times\mathcal{F}\mapsto\mathbb{R}_{\geq 0}$ so that $d(f_{i},f_{j})$ measures the distance (change) between models $i$ and $j$. We remark that the stability metric could be based on parameters (in parametric models), structure (in tree-based models), or even predictions. Then, for each pair of consecutive models, we minimize $\mathcal{L}_{S}(\bm{f})=\sum_{b=1}^{B-1}d(f_{b},f_{b+1})$.
Observation 1.
The stability loss is data-independent, i.e., $\mathcal{L}_{S}(\bm{f},\bm{D})=\mathcal{L}_{S}(\bm{f})$ for all data instances $\bm{D}$, so the stability objective generalizes trivially, i.e., ^$\hat{\mathcal{L}}_{S}(\bm{f})=\mathcal{L}_{S}(\bm{f})$.
@@
2.2 A General Formulation
The objectives stated in Section 2.1 naturally give rise to a multi-objective optimization problem [30]: we search for a sequence $\bm{f}$ that “performs well” in terms of both ^$\hat{\mathcal{L}}_{P}$ and ^$\hat{\mathcal{L}}_{S}$. This is commonly achieved by minimizing (i) a weighted combination of the objectives, i.e., $\lambda\hat{\mathcal{L}}_{P}+(1-\lambda)\hat{\mathcal{L}}_{S}$, where the weight $\lambda\in[0,1]$ is selected based on the learner’s preferences (assuming these are known), or (ii) a preference-agnostic, min-max type formulation [31]. In our setting, we often want to explicitly control “the price of stability” with respect to predictive power. This is particularly relevant in healthcare applications, where the accuracy requirements for using a model in production are stringent. Therefore, we opt for an $\epsilon$-constrained formulation: given a suboptimality tolerance parameter $\alpha\geq 0$, the predictive error of each model $f_{b}$ should not be worse than the best model for batch $D_{b}$ by more than a multiplicative factor of $\alpha$. Recalling our assumption that, at training time, we have already received $B\geq 2$ batches of data (if this is not the case, we artificially split the training data into $B$ batches), we obtain the following formulation:
$\begin{split}\min_{f_{1},\ldots,f_{B}\in\mathcal{F}^{B}}&\quad\hat{\mathcal{L}}_{S}(\bm{f})\ :=\ \sum_{b=1}^{B-1}d(f_{b},f_{b+1})\\ \text{s.t.}&\quad\hat{\mathcal{L}}_{P}(f_{b},D_{b})\leq(1+\alpha)\cdot\min_{f\in\mathcal{F}}\hat{\mathcal{L}}_{P}(f,D_{b}):=\epsilon_{b}\quad\forall b\in[B].\end{split}$ (1)
@@ 2.2 A General Formulation
A sequence $\bm{f}\in\mathcal{F}^{B}$ is weakly Pareto optimal (WPO) if there is no other sequence $\bm{f^{\prime}}$ such that ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})<\hat{\mathcal{L}}_{P}(f_{b},D_{b})\ \forall b\in[B]$ and ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})<\hat{\mathcal{L}}_{S}(\bm{f})$. $\bm{f}$ is Pareto optimal (PO) if the inequality is strict for at least one objective and non-strict for the remaining ones.
@@
The next results show that any solution to Problem (1) is WPO, whereas a unique solution is PO:
Theorem 1.
A sequence $\bm{f^{*}}$ obtained by solving Problem (1) is WPO.
Proof.
(Proof of Theorem 1) Let $\bm{f^{*}}$ be a solution to Problem (1). Assume $\bm{f^{*}}$ is not WPO. Then, $\exists\bm{f^{\prime}}$ such that ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})<\hat{\mathcal{L}}_{P}(f^{*}_{b},D_{b})\leq\epsilon_{b}\ \forall b$ (so $\bm{f^{\prime}}$ is feasible) and ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})<\hat{\mathcal{L}}_{S}(\bm{f^{*}})$. This contradicts the optimality of $\bm{f^{*}}$ among feasible $\bm{f}$; thus, $\nexists$ such $\bm{f^{\prime}}$ so $\bm{f^{*}}$ is WPO. ∎
Theorem 2.
A sequence $\bm{f^{*}}$ is PO if it is a unique solution to Problem (1) for any given $\bm{\epsilon}$.
Proof.
(Proof of Theorem 2) Let $\bm{f^{*}}$ be a unique solution to Problem (1). Then, ^$\hat{\mathcal{L}}_{S}(\bm{f^{*}})<\hat{\mathcal{L}}_{S}(\bm{f})\ \forall\bm{f}$ when ^$\hat{\mathcal{L}}_{P}(f_{b},D_{b})\leq\epsilon_{b}\ \forall b$. Assume $\bm{f^{*}}$ is not PO. Then, $\exists\bm{f^{\prime}}$ such that ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})\leq\hat{\mathcal{L}}_{P}(f^{*}_{b},D_{b})\ \forall b$, ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})\leq\hat{\mathcal{L}}_{S}(\bm{f^{*}})$, and the inequality is strict for at least one $b^{\prime}\in[B]$ or for the stability loss. If the inequality is strict for the stability loss, then we have ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})<\hat{\mathcal{L}}_{S}(\bm{f^{*}})$ (so $\bm{f^{\prime}}$ achieves a better objective value in Problem (1)) and ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})\leq\hat{\mathcal{L}}_{P}(f^{*}_{b},D_{b})\leq\epsilon_{b}\ \forall b$ (so $\bm{f^{\prime}}$ is feasible for Problem (1)). Thus, we have a contradiction with the fact that $\bm{f^{*}}$ solves Problem (1). If the inequality is strict for $b^{\prime}\in[B]$, then ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b^{\prime}},D_{b^{\prime}})<\hat{\mathcal{L}}_{P}(f^{*}_{b^{\prime}},D_{b^{\prime}})\leq\epsilon_{b^{\prime}}$, ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})\leq\hat{\mathcal{L}}_{P}(f^{*}_{b},D_{b})\leq\epsilon_{b}\ \forall b\not=b^{\prime}$, and ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})\leq\hat{\mathcal{L}}_{S}(\bm{f^{*}})$. This contradicts the uniqueness of $\bm{f^{*}}$. Thus, $\nexists$ such $\bm{f^{\prime}}$ and $\bm{f^{*}}$ is PO. ∎
@@
In light of Theorems 1 and 2, the proposed methodology is guaranteed to produce PO sequences for models that correspond to strongly convex optimization problems (e.g., $\ell_{2}$-regularized logistic regression) and WPO sequences for all models. As we show next, to verify PO for non-strongly convex optimization problems, it suffices to solve a series of “complementary problems,” in each of which the predictive power loss for one batch is minimized subject to appropriately chosen constraints on the remaining predictive power losses and the stability loss:
Theorem 3.
@@
A sequence $\bm{f^{*}}$ is PO if and only if it solves both Problem (1) with $\epsilon_{b}=\hat{\mathcal{L}}_{P}(f^{*}_{b},D_{b}):=\epsilon^{*}_{b}\ \forall b\in[B]$ and the following problem for all $b^{\prime}\in[B]$:
@@
$\min_{\bm{f}}\hat{\mathcal{L}}_{P}(f_{b^{\prime}},D_{b^{\prime}})\quad\text{s.t.}\ \hat{\mathcal{L}}_{P}(f_{b},D_{b})\leq\epsilon^{*}_{b}\ \forall b\not=b^{\prime},\quad\hat{\mathcal{L}}_{S}(\bm{f})\leq\hat{\mathcal{L}}_{S}(\bm{f}^{*}).$ (2)
Proof.
(Proof of Theorem 3) Let $\bm{f^{*}}$ be a solution to Problem (1) with $\epsilon_{b}=\epsilon^{*}_{b}\ \forall b$ and the complementary Problem (2) $\forall b^{\prime}\in[B]$. Then, by optimality of $\bm{f^{*}}$ for Problem (1), $\nexists\bm{f^{\prime}}$ such that ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})<\hat{\mathcal{L}}_{S}(\bm{f}^{*})$ and ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})\leq\epsilon^{*}_{b}\ \forall b$. Further, by optimality of $\bm{f^{*}}$ for Problem (2) for all $b^{\prime}\in[B]$, $\nexists\bm{f^{\prime}}$ such that ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b^{\prime}},D_{b^{\prime}})<\epsilon^{*}_{b^{\prime}}$, ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})\leq\hat{\mathcal{L}}_{S}(\bm{f}^{*})$, and ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})\leq\epsilon^{*}_{b}\ \forall b\not=b^{\prime}$. Thus, $\bm{f^{*}}$ is PO since, in both cases, the inequality is strict for at least one objective.
For the opposite direction, let $\bm{f^{*}}$ be PO. First, assume $\bm{f^{*}}$ does not solve Problem (1) with $\epsilon_{b}=\epsilon^{*}_{b}\ \forall b$. Then, $\exists\bm{f^{\prime}}$ that is feasible, i.e., ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})\leq\epsilon^{*}_{b}\ \forall b$, such that ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})<\hat{\mathcal{L}}_{S}(\bm{f}^{*})$. This contradicts the PO of $\bm{f^{*}}$. Now, assume $\bm{f^{*}}$ does not solve Problem (2) for some $b^{\prime}$. Then, $\exists\bm{f^{\prime}}$ that is feasible, i.e., ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})\leq\epsilon^{*}_{b}\ \forall b\not=b^{\prime}$ and ^$\hat{\mathcal{L}}_{S}(\bm{f^{\prime}})\leq\hat{\mathcal{L}}_{S}(\bm{f}^{*})$, such that ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b^{\prime}},D_{b})<\epsilon^{*}_{b^{\prime}}$. This also contradicts the PO of $\bm{f^{*}}$. Thus, $\bm{f^{*}}$ solves the series of Problems stated in Theorem 3. ∎
@@
In all, we can recover the entire (empirical) Pareto frontier by varying the value of the accuracy tolerance parameter in Problem (1), and possibly solving a series of complementary problems for verification. The next result, which follows by direct application of Theorem 4 from [32], guarantees that, by finding the empirical Pareto frontier, we approximately recover the true Pareto frontier with respect to all objectives:
Lemma 4.

### v5

@@ Abstract
We consider the problem of retraining machine learning (ML) models when new batches of data become available. Existing approaches greedily optimize for predictive power independently at each batch, without considering the stability of the model’s structure or analytical insights across retraining iterations. We propose a model-agnostic framework for finding sequences of models that are stable across retraining iterations. We develop a mixed-integer optimization formulation that is guaranteed to recover Pareto optimal models (in terms of the predictive power-stability trade-off) with good generalization properties, as well as an efficient polynomial-time algorithm that performs well in practice. We focus on retaining consistent analytical insights — which is important to model interpretability, ease of implementation, and fostering trust with users — by using custom-defined distance metrics that can be directly incorporated into the optimization problem. We evaluate our framework across models (regression, decision trees, boosted trees, and neural networks) and application domains (healthcare, vision, and language), including deployment in a production pipeline at a major US hospital. We find that, on average, a 2% reduction in predictive power leads to a 30% improvement in stability.
@@ 1 Introduction
Machine learning models deployed in real-world production pipelines often require retraining to incorporate new data as it becomes available Gama et al. (2014); Wu et al. (2020). Retraining involves updating the model periodically to ensure its predictions remain accurate and aligned with the evolving data. However, in high-stakes domains such as healthcare, this process presents challenges. Frequent updates can lead to significant changes in the model’s structure and behavior, undermining trust among expert users Glikson & Woolley (2020) and complicating the model’s integration within operational pipelines Baier et al. (2019).
In sensitive applications, trust in AI-driven decision support systems is critical and is built over time. Physicians, policymakers, and other domain experts rely not only on a model’s accuracy but also on its consistency and alignment with domain knowledge Dietvorst et al. (2018); Orfanoudaki et al. (2022). Retraining that leads to unpredictable shifts in the model’s structure or analytical insights can cause skepticism, hesitation in adoption, or even rejection of the model Longoni et al. (2019); Babic et al. (2021). Moreover, unstable retraining processes may result in legal or operational complications due to errors in critical decisions Bertsimas & Orfanoudaki (2021).
Stability Turney (1995); Bousquet & Elisseeff (2002) plays a pivotal role in overcoming these challenges. Stability in model structures ensures that changes between retraining iterations are smooth and interpretable, preserving trust in the system while allowing users to incorporate new insights without losing historical context. This is particularly important for high-stakes applications where consistent analytical insights are as valuable as predictive performance.
1.1 Contributions
We propose a framework to explicitly enforce stability in model retraining. Unlike existing methods, which focus primarily on maximizing predictive performance, our approach balances predictive power and structural stability. Our contributions are as follows:
  • Framework for Retraining: We propose a general, model-agnostic framework for retraining ML models that explicitly enforces stability across retraining iterations. The framework is also data-agnostic:
    • If the data distribution remains stable over time, our framework prioritizes sequences of models that exhibit minimal structural changes, maintaining consistent insights.
    • If the data distribution shifts significantly, our framework naturally favors models that better align with the new data, while still minimizing abrupt structural changes that undermine trust.
  • Theoretical Guarantees: We provide rigorous guarantees, showing that our framework produces solutions that are Pareto optimal or weakly Pareto optimal with respect to predictive performance and stability. We also establish a generalization bound for our methodology.
  • Algorithm for Practical Implementation: Leveraging a mixed-integer optimization (MIO) formulation, we develop an efficient heuristic algorithm to make our framework practical for real-world applications. The algorithm accommodates a variety of model types, including logistic regression, classification trees, boosted trees, and neural networks.
  • Numerical Validation: We validate our methodology across three real-world case studies, demonstrating its ability to improve stability with a small, controllable trade-off in predictive performance. This includes deployment in a production pipeline at a large US hospital, where our approach ensures stable mortality prediction models in a high-stakes healthcare setting.
1.2 Related Work
Batch vs. Online Learning. Retraining ML models is a common requirement in production pipelines to incorporate new data and maintain predictive performance Gama et al. (2014); Wu et al. (2020). Depending on the data update frequency, the ML model retraining problem can be viewed either as a batch or an online learning problem. Batch learning involves periodically retraining models from scratch using the full dataset, whereas online learning incrementally updates models as new data becomes available Bisong (2019); Hoi et al. (2021). Online learning methods are computationally efficient but typically rely on partial data, resulting in greedier solutions that may not generalize as well. Recent works in online learning and adaptive learning Gama et al. (2014) highlight their utility in dynamic environments, but these approaches often ignore structural consistency and require frequent updates that can disrupt production pipelines. In contrast, batch learning, by leveraging the entire dataset for retraining, ensures globally optimal solutions and smoother updates, making it better suited for high-stakes domains like healthcare, where interpretability, trust, and seamless integration with existing operational pipelines are critical.
Model Retraining Challenges: Distribution Shifts and Costs. Retraining ML models often requires addressing challenges such as detecting distribution shifts, managing retraining costs, and ensuring smooth transitions. Existing works focus on detecting shifts Bifet & Gavalda (2007); Pesaranghader & Viktor (2016) or adapting models to mitigate their effects Schwinn et al. (2022); as shown by Kabra & Patel (2024), in settings where the deployed model influences the data distribution, naive retraining strategies can largely fail. Several works also consider retraining costs: Mahadevan & Mathioudakis (2023) propose an algorithm that balances retraining costs with performance deterioration by approximating a “staleness” cost, whereas Žliobaitė et al. (2015) propose a return-on-investment framework for assessing retraining needs. However, these approaches primarily focus on deciding when to retrain rather than how to ensure consistency across retraining iterations, possibly resulting in significant changes to model behavior. Our work bridges these gaps by explicitly incorporating structural stability into the retraining process, enabling smooth transitions even under significant distribution shifts.
Stable and Interpretable Models. The importance of stability for trust and interpretability has been recognized since early works Turney (1995); Breiman (1996). Recent methods address stability either by training interpretable models with inherent structural consistency Dwyer & Holte (2007); Bargagli-Stoffi et al. (2020); Bertsimas & Digalakis Jr (2023); Bertsimas et al. (2024) or by ensuring stable feature interactions in black-box models Basu et al. (2018); Wang et al. (2024). Beyond classical ML models, Zhang et al. (2020) propose a retraining approach for recommender systems that, upon the arrival of new data, transfers past knowledge to the new model: while their method prioritizes efficiency and accuracy, it enforces an indirect form of stability by transferring learned experiences. These approaches are limited to specific model types or focus solely on feature importance without considering the stability of entire model sequences. Our work complements this by explicitly prioritizing structural stability alongside predictive performance, offering a unified framework that generalizes across model types and applications.
2 A Framework for ML Model Retraining
@@ 2.1 Problem Setting
Consider an observational dataset accumulated over time composed of triplets $(\boldsymbol{x_{i}},y_{i},t_{i})_{i=1}^{N}$, where $\boldsymbol{x_{i}}\in\mathcal{X}\subseteq\mathbb{R}^{p}$ represents the feature vector, $y_{i}\in\mathcal{Y}\subseteq\mathbb{R}$ is the target variable, and $t_{i}\in\mathbb{R}$ is the time at which data point $i$ was obtained. The dataset is divided into a series of batches at times $T_{1}<T_{2}<\ldots<T_{B}$. The dataset available at time $T_{b},\ b\in[B]:=\{1,\ldots,B\},$ is defined as $D_{b}=\{(x_{i},y_{i}):t_{i}\leq T_{b},i\in[N]\}.$ Thus, we have that $D_{1}\subseteq D_{2}\subseteq\ldots\subseteq D_{B}$ and $\boldsymbol{D}=(D_{1},\dots,D_{B})$. We assume that $B\geq 2$ and that, at model training time (denoted by $t$), all $B$ batches are available, i.e., $t\geq T_{B}$. In Section 3.2, we discuss how to adaptively retrain the model using batches obtained after time $t$. The goal is to find models $\boldsymbol{f}:=(f_{1},f_{2},\ldots,f_{B})\in\mathcal{F}\times...\times\mathcal{F}:=\mathcal{F}^{B}$ with $f_{b}:\mathcal{X}\mapsto\mathcal{Y}$ for $b\in[B]$ such that the following two objectives are satisfied:
Predictive power: The first objective is to ensure that each model $f_{b}$ in the sequence has small expected loss in terms of prediction quality. We denote by $\mathbb{P}(D_{b})$ the distribution of the training data in $D_{b}$ and $\ell:\mathcal{Y}\times\mathcal{Y}\mapsto\mathbb{R}_{\geq 0}$ the bounded loss function measuring, e.g., regression or classification accuracy; concretely, $\ell(y,f_{b}(\boldsymbol{x}))$ quantifies the quality of model $f_{b}$’s prediction on input $\boldsymbol{x}$ for $y$. Then, we seek for a sequence whereby each model (individually) achieves low expected loss $\mathcal{L}_{P}(f_{b},D_{b})=\mathbb{E}_{(\boldsymbol{x},y)\sim\mathbb{P}(D_{b})}[\ell(y,f_{b}(\boldsymbol{x}))].$ The overall loss of the sequence is then denoted by $\mathcal{L}_{P}(\boldsymbol{f},\boldsymbol{D})=\sum_{b=1}^{B}\mathcal{L}_{P}(f_{b},D_{b}).$ In practice, we approximate the expected loss using the empirical loss ^$\hat{\mathcal{L}}_{P}(f_{b},D_{b})=\frac{1}{N_{b}}\sum_{i:t_{i}\leq T_{b}}[\ell(y_{i},f_{b}(\boldsymbol{x}_{i}))]$. The next assumption bounds the approximation quality as function of the sample size, the function class complexity, and the desired confidence:
Assumption 2.1.
@@
Stability: The second objective is to ensure that consecutive models do not vary significantly, maintaining stability over time. Formally, we define a distance metric $d:\mathcal{F}\times\mathcal{F}\mapsto\mathbb{R}_{\geq 0}$ so that $d(f_{i},f_{j})$ measures the change between models $i$ and $j$. The distance metric can be based on the models’ parameters or structure (e.g., for linear regressions $\boldsymbol{\beta}_{i},\boldsymbol{\beta}_{j}$, we can simply take $d(\boldsymbol{\beta}_{i},\boldsymbol{\beta}_{j})=\|\boldsymbol{\beta}_{i}-\boldsymbol{\beta}_{j}\|^{2}$), analytical insights (e.g., feature importances), or even predictions; we discuss the choice of distance metric in detail in Section 4. Then, for each pair of consecutive models, we minimize the stability loss $\mathcal{L}_{S}(\boldsymbol{f})=\sum_{b=1}^{B-1}d(f_{b},f_{b+1})$.
Remark 2.2.
The stability loss is data-independent, i.e., $\mathcal{L}_{S}(\boldsymbol{f},\boldsymbol{D})=\mathcal{L}_{S}(\boldsymbol{f})$ for all data instances $\boldsymbol{D}$, so the stability objective generalizes trivially, i.e., ^$\hat{\mathcal{L}}_{S}(\boldsymbol{f})=\mathcal{L}_{S}(\boldsymbol{f})$.
@@
2.2 General Formulation
The objectives stated in Section 2.1 naturally give rise to a multi-objective optimization problem Miettinen (1999): we search for a sequence $\boldsymbol{f}$ that “performs well” in terms of both ^$\hat{\mathcal{L}}_{P}$ and ^$\hat{\mathcal{L}}_{S}$. This is commonly achieved by minimizing (i) a weighted combination of the objectives, i.e., $\lambda\hat{\mathcal{L}}_{P}+(1-\lambda)\hat{\mathcal{L}}_{S}$, where the weight $\lambda\in[0,1]$ is selected based on the learner’s preferences (assuming these are known), or (ii) a preference-agnostic, min-max type formulation Cortes et al. (2020). In our setting, we often want to explicitly control “the price of stability” with respect to predictive power. This is particularly relevant in healthcare applications, where the accuracy requirements for using a model in production are stringent. Therefore, we opt for an $\epsilon$-constrained formulation: given a suboptimality tolerance parameter $\alpha\geq 0$, the predictive error of each model $f_{b}$ should not be worse than the best model for batch $D_{b}$ by more than a multiplicative factor of $\alpha$. Recalling our assumption that, at training time, we have already received $B\geq 2$ batches of data (if this is not the case, we artificially split the training data into $B$ batches), we obtain the following formulation:
$\begin{split}\min_{f_{1},\ldots,f_{B}\in\mathcal{F}^{B}}&\quad\hat{\mathcal{L}}_{S}(\boldsymbol{f})\ :=\ \sum_{b=1}^{B-1}d(f_{b},f_{b+1})\\ \text{s.t.}&\quad\hat{\mathcal{L}}_{P}(f_{b},D_{b})\leq(1+\alpha)\cdot\\ &\quad\cdot\min_{f\in\mathcal{F}}\hat{\mathcal{L}}_{P}(f,D_{b}):=\epsilon_{b}\quad\forall b\in[B].\end{split}$ (1)
2.3 Pareto Optimality of Solutions
@@
A sequence $\boldsymbol{f}\in\mathcal{F}^{B}$ is weakly Pareto optimal (WPO) if there is no other sequence $\boldsymbol{f^{\prime}}$ such that ^$\hat{\mathcal{L}}_{P}(f^{\prime}_{b},D_{b})<\hat{\mathcal{L}}_{P}(f_{b},D_{b})\ \forall b\in[B]$ and ^$\hat{\mathcal{L}}_{S}(\boldsymbol{f^{\prime}})<\hat{\mathcal{L}}_{S}(\boldsymbol{f})$. $\boldsymbol{f}$ is Pareto optimal (PO) if the inequality is strict for at least one objective and non-strict for the remaining ones.
@@
The next results, which we prove in Appendix A, show that any solution to Problem (1) is WPO, whereas a unique solution is PO:
Theorem 2.4.
A sequence $\boldsymbol{f^{*}}$ obtained by solving Problem (1) is WPO.
Theorem 2.5.
A sequence $\boldsymbol{f^{*}}$ is PO if it is a unique solution to Problem (1) for any given $\boldsymbol{\epsilon}$.
@@
In light of Theorems 2.4 and 2.5, the proposed methodology is guaranteed to produce PO sequences for models that correspond to strongly convex optimization problems (e.g., $\ell_{2}$-regularized logistic regression) and WPO sequences for all models. As we show next (proof in Appendix A), to verify PO for non-strongly convex optimization problems, it suffices to solve a series of “complementary problems,” in each of which the predictive power loss for one batch is minimized subject to appropriately chosen constraints on the remaining predictive power losses and the stability loss:
Theorem 2.6.
@@
A sequence $\boldsymbol{f^{*}}$ is PO if and only if it solves both Problem (1) with $\epsilon_{b}=\hat{\mathcal{L}}_{P}(f^{*}_{b},D_{b}):=\epsilon^{*}_{b}\ \forall b\in[B]$ and the following problem for all $b^{\prime}\in[B]$:
@@
$\begin{split}\min_{\boldsymbol{f}}&\quad\hat{\mathcal{L}}_{P}(f_{b^{\prime}},D_{b^{\prime}})\\ \text{s.t.}&\quad\hat{\mathcal{L}}_{P}(f_{b},D_{b})\leq\epsilon^{*}_{b}\quad\forall b\not=b^{\prime}\\ &\quad\hat{\mathcal{L}}_{S}(\boldsymbol{f})\leq\hat{\mathcal{L}}_{S}(\boldsymbol{f}^{*}).\end{split}$ (2)
@@
In all, we can recover the entire (empirical) Pareto frontier by varying the value of the accuracy tolerance parameter in Problem (1), and possibly solving a series of complementary problems for verification.
2.4 Pareto Excess Risk Bound
The next result, which follows by direct application of Theorem 4 from Súkeník & Lampert (2024), guarantees that, by finding the empirical (in-sample) Pareto frontier, we approximately recover the true (out-of-sample) Pareto frontier with respect to all objectives:
Lemma 2.7.